OpenStudy (anonymous):

Solve the following system of equations for the unknown variables. y = 2x x = 2y – 2

OpenStudy (anonymous):

sutitution x = 2y – 2

OpenStudy (anonymous):

wait let me check

OpenStudy (anonymous):

yeah last answer came up wrong would u like possible answer's?

OpenStudy (anonymous):

y=2x sutitution x=2y -2 y=2(2y-2) y=4y-4 4y -y = 4 3y= 4 y = 4/3

OpenStudy (anonymous):

now sutitution y =4/3 solve for x x =2(4/3) -2 x=(8/3)-2 x=(8-(2*3))/3 x =(8-6)/3 x= 2/3

OpenStudy (anonymous):

here is correct answer

OpenStudy (anonymous):

yeah thats the one thank u how do you know when to substitute -2?

OpenStudy (anonymous):

what u mean, the frist I solve is correct equation, but when I caculator I was wrong

OpenStudy (anonymous):

-2 form this equation x = 2y – 2

OpenStudy (anonymous):

im having issues with where ur getting the -2

OpenStudy (anonymous):

r u getting now?

OpenStudy (anonymous):

i understand the process of getting the answer but do i always use neg two or will it change with different equations?

OpenStudy (anonymous):

frist use ; y=2x

OpenStudy (anonymous):

and sutitution x = 2y – 2

OpenStudy (anonymous):

y=2( ) plug 2y – 2 inside Parentheses and solve

OpenStudy (anonymous):

after find y =4/3 use eqution x = 2y – 2 x = 2( ) – 2 plug 4/3 inside Parentheses and solve

OpenStudy (anonymous):

2 way can solve * use subtitution *elemate this given value x and y you use sutitution more easy

OpenStudy (anonymous):

know what u mean 2x-y=0 -x+2y=2 (this equation time 2) u get -2x+4y=4 -2x+4y=4 Add 2x-y=0 ----------------------- 3y=4 y=4/3 this u mean? answer come up the same

OpenStudy (anonymous):

ahh ok see i was confused which method to use substitution seems easier

OpenStudy (anonymous):

if u have value x and y better do sutitution more easy if standart from use elemate

OpenStudy (anonymous):

cool ive got another one im haven issues with u wana check it out

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Solve the following system of equations for the unknown variables. 3x + 3y = 3 -3x – 2z = -11 y + z = -1

OpenStudy (anonymous):

solving for x,y,z

OpenStudy (anonymous):

u know how to get rid by negative side?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

this standart from than solve by elemate

OpenStudy (anonymous):

give me time this one is hard

OpenStudy (anonymous):

yeah i figured out the last two but this one i keep running into problems

OpenStudy (anonymous):

Sorry u need u sutitution not elemate y + z = -1 y= -1-z 3x + 3y = 3 (plug y= -1-z in this eqution u get) 3x + 3(-1-z) = 3

OpenStudy (anonymous):

can u solve next?

OpenStudy (anonymous):

yeah i think so if z=-1 i just plug it into the other equations to get x and y

OpenStudy (anonymous):

yes, I think u can do the rest

OpenStudy (anonymous):

thanks how did u get so good at math

OpenStudy (anonymous):

y + z = -1 (subtitution z= -1) y+(-1) =-1 y -1 = -1 y =-1+1 y= 0

OpenStudy (anonymous):

yeah i got x=1 y=0 z=-1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

makes sense i think i messed up when i put y=-z-1

OpenStudy (anonymous):

let me check, some time I caculator wrong

OpenStudy (anonymous):

z=1 not -1 check caculator

OpenStudy (anonymous):

3x-3z=6 -3x – 2z = -11 -------------------------------- Add -5z=-5 z=1

OpenStudy (anonymous):

thats gana change my other answers ill have to go back through

OpenStudy (anonymous):

y + z = -1 (subtitution z= 1) y+1 =-1 y = -1-1 y =-2

OpenStudy (anonymous):

3x + 3y = 3 3x + 3(-2) = 3 3x-6=3 3x=3+6 x=9/3 x=3

OpenStudy (anonymous):

i got same for x

OpenStudy (anonymous):

good now plug and check

OpenStudy (anonymous):

I plug z=1 x=3 y=-2 the answer is correct

OpenStudy (anonymous):

i got same those equations take forever!

OpenStudy (anonymous):

I'm glad u can solve problem now ,some time u have to use sutittution

OpenStudy (anonymous):

i can do the simple ones no problem but when they get complex i tend to get mixed up in my steps

OpenStudy (anonymous):

this problem frist step u have to use sutitution

OpenStudy (anonymous):

u know now?

OpenStudy (anonymous):

yeah i was using the same method just got mixed up in the initial steps for finding z

OpenStudy (anonymous):

I tutor my son he like solve by sbutitution instate elemate

OpenStudy (anonymous):

cool im doing an electricians math course its been a very long time since ive done systems of equations

OpenStudy (anonymous):

ok bye

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np